1. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many NET ATP are generated in these reactions? ("Net" refers to total # generated - total number consumed.)

2. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many CO2 molecules are released in these reactions? 3. If one glucose molecule enters into glycolysis and proceeds through the PDC to acetyl-CoA, how many NADH molecules are generated in these reactions?
Answer #1

1. Glycolysis -

4 ATP synthesized

2 ATP consumed

2 NADH synthesized (6 ATP)

Pyruvate decarboxylation -

2 NADH synthesized (6 ATP)

ATP synthesized = 4 + 6 + 6 = 16

ATP consumed = 2

Net ATP generated = 16 - 2 = 14

2. Two only

0 carbondioxide consumed/produced in glycolysis

1 carbondioxide produced in pyruvate decarboxylation. This occurs twice for one molecule of glucose.

3. 4 NADH (explained in 1)

Please rate high.